OpenStudy (anonymous):

Slope-intercept form: 1. parallel to y=4x+2, y-intercept at 4 2. perpendicular to y=1/4x+2, passes through (0,0) 3. slope -2/3, x-intercept at 3

jimthompson5910 (jim_thompson5910):

all parallel lines have equal slopes

jimthompson5910 (jim_thompson5910):

so anything parallel to y = 4x+2 has a slope of 4

OpenStudy (anonymous):

ok so how do i solve the first one?

jimthompson5910 (jim_thompson5910):

we have a slope of 4 and a y-intercept of 4

jimthompson5910 (jim_thompson5910):

general equation is y = mx+b

jimthompson5910 (jim_thompson5910):

m = slope, b = y-intercept

OpenStudy (anonymous):

ok i got that part

OpenStudy (anonymous):

so would it be 4=4x+2?

jimthompson5910 (jim_thompson5910):

no, y = 4x+4

OpenStudy (anonymous):

ok then don't you subtract 4 from both sides to b or leave as it is?

jimthompson5910 (jim_thompson5910):

you leave it as is

OpenStudy (anonymous):

ok what about problem 2?

jimthompson5910 (jim_thompson5910):

Take the slope of 1/4 and flip it to get 4/1 or 4. Then change the sign to get -4

jimthompson5910 (jim_thompson5910):

So y = mx+b y = -4x+b ... Plug in m = -4 0 = (-4)(0) + b Now solve for b

OpenStudy (anonymous):

well -4 times 0 equals 0 so what would b equal?

jimthompson5910 (jim_thompson5910):

go so far y = mx+b y = -4x+b ... Plug in m = -4 0 = (-4)(0) + b 0 = 0 + b 0 = b b = 0

jimthompson5910 (jim_thompson5910):

good*

jimthompson5910 (jim_thompson5910):

So the equation is y = -4x+0 or just y = -4x

OpenStudy (anonymous):

ok thanks what about 3?

jimthompson5910 (jim_thompson5910):

y = mx + b y = (-2/3)x + b ... Plug in m = -2/3 0 = (-2/3)(3) + b ... Now plug in the given x-intercept 0 = -6/3 + b 0 = -2 + b 0+2 = -2 + b + 2 2 = b b = 2 So the equation is \(\Large y = -\frac{2}{3}x+2\)

OpenStudy (anonymous):

ok thx

jimthompson5910 (jim_thompson5910):

np